OpenStudy (anonymous):

write the standard form of the equation x^2-6x+y^2-10y-2=0. find the center a. (-3,-5) b.(9,25) c.(-9,-25) d.(3,5)

OpenStudy (lgbasallote):

put 2 on the other side \[x^2 - 6x + y^2 - 10y = 2\] \[(x^2 - 6x + 9) + (y^2 - 10y + 25) = 2 + 9 + 25\] \[(x-3)^2 + (y-5)^2 = 36\] doe that make sense?

OpenStudy (anonymous):

Completion Of Square Method..

OpenStudy (anonymous):

got it. so the center is (3,5)

OpenStudy (lgbasallote):

yes :D

OpenStudy (anonymous):

THANKS!!!!!!! :D

OpenStudy (lgbasallote):

<tips hat>

OpenStudy (anonymous):

do you know how to solve this?

OpenStudy (anonymous):

16x2 + 25y2 - 96x - 200y + 144 = 0. Write the standard form of the ellipse. 1) Find the center 2) Find the foci 3) Find the vertices

OpenStudy (lgbasallote):

sorry idk abt ellipses try asking in a new post

OpenStudy (anonymous):

ok no problem thanks!